OpenStudy (anonymous):

If a function y = f(x) is di ffrentiable at x = 3 and f '(3) = 5, and find the limit lim ((x^(2)-3x))/(f(3)-f(x)) x goes to 3

OpenStudy (kainui):

If you try to plug in a 3 you find that you get 0/0 which is an indeterminate form. So you can use L'hopital's rule which says that the derivative of the top divided by the derivative of the bottom is equal to the limit.

OpenStudy (anonymous):

thank you for your response. It seems if I use that method, then the information given in the question where f prime 3 is equal to 5 is not useful

OpenStudy (anonymous):

do you think Lhopital rule is the only method for this question?

OpenStudy (anonymous):

thank you for the graph. I fully understand what you mean

OpenStudy (kainui):

Actually I don't know, maybe.

OpenStudy (kainui):

But by L'hopital's rule I got the limit as x->3 to be infinity.

OpenStudy (zarkon):

\[\frac{d}{dx}f(3)=0\]

OpenStudy (anonymous):

yep I got infinity as well I got zero for the denominator

OpenStudy (kainui):

d/dx(f(3))=5 Zarkon

OpenStudy (zarkon):

\[\frac{d}{dx}[f(3)-f(x)]=-f'(x)\]

OpenStudy (kainui):

Reread the problem, it's stated.

OpenStudy (zarkon):

no...\[\left.\frac{d}{dx}f(x)\right|_{3}=5\]

OpenStudy (zarkon):

f(3) is just a number the derivative of a number is zero

OpenStudy (anonymous):

did you get 3/-5 as the answer Zarkon?

OpenStudy (zarkon):

yes

OpenStudy (kainui):

"y = f(x) is diffrentiable at x = 3 and f '(3) = 5"

OpenStudy (kainui):

f'(3)=d/dx(f(3))

OpenStudy (zarkon):

that is incorrect

OpenStudy (kainui):

Yeah, I you're right.

OpenStudy (anonymous):

((x^(2)-3x))/(f(3)-f(x))= -x/ ((f(x)-f(3))/(x-3)) lim (f(x)-f(3))/(x-3) x goes to 3 = f'(3)=5 so the answer is -3/5